OpenStudy (anonymous):

Please help, I've been trying to figure this out for half an hour? Integral substitution

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\frac{ x^2 }{ \sqrt{2x+1} }\]

OpenStudy (anonymous):

Whatever I do, I am not getting the right answer!!

OpenStudy (anonymous):

I set \[u=2x+1\] and got \[\frac{ du }{ dx }=2\] and then \[\frac{ 1 }{ 2 }du=dx\]

OpenStudy (anonymous):

Can someone help guide me through the next few steps? I can't figure out what I'm doing wrong

OpenStudy (zarkon):

\[x=\frac{u-1}{2}\]

OpenStudy (zarkon):

\[x^2=\left(\frac{u-1}{2}\right)^2=\frac{1}{4}(u^2-2u+1)\]

OpenStudy (anonymous):

How did you get u-1?

OpenStudy (zarkon):

solve \(u=2x+1\) for x

OpenStudy (anonymous):

Is this correct: \[\frac{ 1 }{ 2 } \ \int\limits_{1}^{2} \frac{ u^2-2u+1 }{ 4 } \times \frac{ u^{^{1/2}} }{ \frac{ 1 }{ 2 } }\]

OpenStudy (zarkon):

why is \(u^{1/2}\) in the numerator?

OpenStudy (anonymous):

Oops. I forgot to take the integral sign away. Does it make sense if I take the sign away? And then would I just evaluate it at x=2 and x-1?

OpenStudy (anonymous):

=*

OpenStudy (zarkon):

you need to integrate first

OpenStudy (zarkon):

the entire integrand

OpenStudy (anonymous):

I did \[\int\limits_{}^{}u ^{-1/2} = \frac{ u^{1/2} }{ \frac{ 1 }{ 2 } }\]

OpenStudy (zarkon):

you can't just integrate that part

OpenStudy (zarkon):

you have to integrate the entire thing

OpenStudy (anonymous):

So I would just integrate everything after the integrate I set up right?

OpenStudy (zarkon):

distribute the \(u^{-1/2}\) then integrate

OpenStudy (anonymous):

integrand sign*

OpenStudy (zarkon):

also you need to change your limits of integration

OpenStudy (anonymous):

Why would I change those?

OpenStudy (zarkon):

1 to 2 is for x you are integrating over u

OpenStudy (anonymous):

So how would I figure out what my new limits of integration are?

OpenStudy (zarkon):

use u=2x+1

OpenStudy (zarkon):

when x=1 u=2(1)+1=3 when x=2 u=2(2)+1=5

OpenStudy (zarkon):

so if 1<x<2 then 3<u<5

OpenStudy (anonymous):

okay thanks